OpenStudy (anonymous):

The Atlanta Hawks scored 92 points in a basketball game. They made a total of 41 baskets. Some baskets were 2-points and some were 3points. Create a system of equations to represent the problem and find the number of 2- and 3-point baskets made by graphing. a)name the variables: x is the number of 2-point baskets and y is . b)set up equation for the number of baskets made: c)set up equation for the number of points scored:

OpenStudy (math&ing001):

41=x+y 92=2x+3y

OpenStudy (math&ing001):

The answer to part a is already given. The question is to name the variables so: a) x is the number of 2-point baskets and y is the number of 3-point baskets. b) 41=x+y c) 92=2x+3y

OpenStudy (anonymous):

Here is the next part: d) Graph the first equation (number of baskets made. You can do this by hand by picking a value for x (number of 2-point baskets) and solving for y. Let’s pick 2 for x. x = 2, y = Therefore, the point (2, ) is on the graph. Let’s pick another value for x to find a second point on the graph. Let x be 4. x = 4, y = Therefore, the point (4, ) is on the graph. Let’s pick one more value for x. Let x be 6. x = 6, y = Therefore, the point (6, ) is on the graph. Graph these three points and sketch the line through them.

OpenStudy (anonymous):

And the last part is: e) Graph the second equation. You can do this by hand by picking a value for x (number of 2-point baskets) and solving for y. Let’s pick 1 for x. x = 1, y = Therefore, the point (1, ) is on the graph. Let’s pick another value for x to find a second point on the graph. Let x be 4. x = 44, y = Therefore, the point (4, ) is on the graph. Let’s pick one more value for x. Let x be 7. x = 7, y = Therefore, the point (7, ) is on the graph. On the same coordinate plane as before, plot these three points and sketch the line through them. f) Locate the point were the two graphs cross. This occurs at ( , ), which means that the number of 2-point baskets (x) scored is , and the number of 3-point baskets (y) scored is .

OpenStudy (math&ing001):

For question d): you plug in the value of x given, in the equation you got in b) (aka 41=x+y) and you find y. For question e): same as d) except you use the second equation 92=2x+3y When you graph the 2 equations they'll cross at a certain, which will be the solution to your system.

OpenStudy (anonymous):

I got it! Thank you so much.